# DMin: Scalable Training Data Influence Estimation for Diffusion Models

### v3

@@ Abstract
Identifying the training data samples that most influence a generated image is a critical task in understanding diffusion models (DMs), yet existing influence estimation methods are constrained to small-scale or LoRA-tuned models due to computational limitations. To address this challenge, we propose DMin (Diffusion Model influence), a scalable framework for estimating the influence of each training data sample on a given generated image. To the best of our knowledge, it is the first method capable of influence estimation for DMs with billions of parameters. Leveraging efficient gradient compression, DMin reduces storage requirements from hundreds of TBs to mere MBs or even KBs, and retrieves the top-$k$ most influential training samples in under 1 second, all while maintaining performance. Our empirical results demonstrate DMin is both effective in identifying influential training samples and efficient in terms of computational and storage requirements.
@@ 1 Introduction
Diffusion models have emerged as powerful generative models, capable of producing high-quality images and media across various applications [8, 36, 23, 38]. Despite their impressive performance, the extremely large scale and complexity of the datasets used for training are often sourced broadly from the internet [33, 35, 34]. This vast dataset diversity allows diffusion models to generate an extensive range of content, enhancing their versatility and adaptability across multiple domains [19, 5]. However, it also means that these models may inadvertently generate unexpected or even harmful content, reflecting biases or inaccuracies present in the training data.
@@
Recently, many studies have explored influence estimation in diffusion models [26, 18, 26, 27, 11]. These methods assign an influence score to each training data sample relative to a generated image, quantifying the extent to which each sample impacts the generation process. For instance, DataInf [18] and K-FAC [26] are influence approximation techniques tailored for diffusion models. However, they are both second-order methods that require the inversion of the Hessian matrix. To approximate this inversion, they must load all the gradients of training data samples across several predefined timesteps. Notably, in the case of the full-precision Stable Diffusion 3 medium model [10], the gradient of entire model requires approximately 8 GB of storage. Collecting gradients for one training sample over 10 timesteps would consume $8\times 10=80$ GB. Scaling this requirement to a training dataset of $10,000$ samples results in a storage demand of around 800 TB – far exceeding the capacity of typical memory or even hard drives. Given that diffusion models are often trained on datasets with millions of samples, this storage demand becomes impractical. Consequently, these methods are limited to LoRA-tuned models or small diffusion models [15, 31]. Although some prior works have applied gradient compression, such as SVD [13] and quantization [26], the achieved compression rates are insufficient to maintain performance at this scale.
@@
Alternatively, Journey-TRAK [11] and D-TRAK [27] are first-order methods for influence estimation on diffusion models, which are extended from TRAK [28] on deep learning models. Both approaches utilize random projection to reduce the dimensionality of gradients. However, for large diffusion models, such as the full-precision Stable Diffusion 3 Medium model, the gradient dimensionality exceeds 2 billion parameters. Using the suggested projection dimension of $32,768$ in D-TRAK, store a such $2\text{B}\times 32,768$ projection matrix requires more than 238 TB of storage. Even projection matrix is dynamically generated during computation, the scale of these operations substantially slows down the overall process. As a result, they are only feasible for small models or adapter-tuned models.
@@
Challenges. Although these approaches have demonstrated superior performance on certain diffusion models, there are several key challenges remain: (1) Scalability on Model Size: Existing methods either require computing second-order Hessian inversion or handling a massive projection matrix, both of which restrict their applicability to large diffusion models. (2) Scalability on Dataset Size: Diffusion models frequently rely on datasets containing millions of samples, making the computation of a Hessian inversion for the entire training dataset impractical. Additionally, storing the full gradients for all training data samples presents a significant challenge. (3) Fragility of Influence Estimation: Previous studies have demonstrated that the fragility of influence estimation in extremely deep models [22, 2, 9, 12]. Similarly, we observed this fragility in large diffusion models, regardless of whether they use U-Net or transformer.
@@
  • Our experimental results demonstrates DMin’s effectiveness and efficiency in influence estimation.
  • We provide an open-source PyTorch implementation with multiprocessing support [^1] https://github.com/huawei-lin/DMin.
@@ 2 Influence Estimation for Diffusion Models
For a latent diffusion model, data $x_{0}$ is first encoded into a latent representation $z_{0}$ using an encoder $E$ by $z_{0}=E(x_{0})$. The model then operates on $z_{0}$ through a diffusion process to introduce Gaussian noise and iteratively denoise it. The objective is to learn to reconstruct $z_{0}$ from a noisy latent $z_{t}$ at any timestep $t\in\{1,2,\cdots,T\}$ in the diffusion process, where $T$ is the number of diffusion steps. Let $\epsilon_{t}\sim\mathcal{N}(0,I)$ denotes the Gaussian noise added at timestep $t$. We define the training objective at each timestep $t$ as follow:
@@
where $\theta$ represents the model parameters, $z_{t}$ is the noisy latent representation of $z_{0}$ at timestep $t$, $f_{\theta}(z_{t},t)$ represent the model’s predicted noise at timestep $t$ for the noisy latent $z_{t}$. $\mathcal{L}(\cdot)$ is the loss function between the predicted noise and actual Gaussian noise.
Given a test generation $x^{s}$, where $x^{s}$ is generated by a well-trained diffusion model with parameters $\theta^{*}$, the goal of influence estimation is to estimate the influence of each training data sample $x^{i}(1\leq i\leq N)$ on generating $s$, where $N$ is the size of training dataset. Let $z^{i}_{t}$ represents the latent representation of $x^{i}$ on timestep $t$, $z^{s}_{t}$ denotes the latent representation of test generation $s$ on timestep $t$.
Considering in the $\alpha$-th training iteration, the model parameter $\theta_{\alpha+1}$ are updated from $\theta_{\alpha}$ by gradient descent on the noise prediction loss for batch $B=(B_{z},B_{t},B_{\epsilon})$:
@@
where $\eta_{\alpha}$ denotes the learning rate in the $\alpha$-th iteration, $(z_{t}^{i},t^{i},\epsilon_{t}^{i})\in B$, and the contribution of $(z_{t}^{i},t^{i},\epsilon_{t}^{i})$ to batch gradient is $\frac{1}{|B|}\nabla_{\theta_{\alpha}}\mathcal{L}(f_{\theta_{\alpha}}(z^{i}_{t},t^{i}),\epsilon^{i}_{t})$. The influence of this training iteration $z^{i}_{t}$ with respect to $z^{s}_{t}$ on timestep $t$ can be quantified as the change in loss:
@@
where $z^{s}_{t}=E(x^{s})+\epsilon^{i}_{t}$, denoting the latent representation of $s$ adding the Gaussian noise $\epsilon^{i}_{t}$, and $\mathcal{I}_{\theta_{\alpha+1},t}(x^{s},x^{i})$ represents the influence of $x^{i}$ with respect to $s$ at the $\alpha$-th iteration with timestep $t$. Then $\mathcal{L}\Bigl{(}f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}$ can be expanded by Taylor expansion:
$\mathcal{L}\Bigl{(}f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}=\mathcal{L}\Bigl{(}f_{\theta_{\alpha}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}$ (4)
$\hskip 14.45377pt+(\theta_{\alpha+1}-\theta_{\alpha})\nabla_{\theta_{\alpha}}\mathcal{L}\Bigl{(}f_{\theta_{\alpha}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}+O(||\theta_{\alpha+1}-\theta_{\alpha}||^{2})$
Given the small magnitude of the learning rate $\eta$, we disregard the higher-order term $O(||\theta_{\alpha+1}-\theta_{\alpha}||^{2})$, as it scales with $O(||\eta_{\alpha}||^{2})$ and is therefore negligible. Then we have:
$\mathcal{I}_{\theta_{\alpha+1},t}(x^{s},x^{i})$ $=\mathcal{L}\Bigl{(}f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}-\mathcal{L}\Bigl{(}f_{\theta_{\alpha}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr{)}$ (5)
@@
To estimate the influence of training data sample $x^{i}$, we can summing up all the training iteration training on $x^{i}$ and timesteps on $t\in\{1,2,\cdots,T\}$:
$\mathcal{I}_{\theta^{*}}(x^{s},x^{i})=$ (6)
@@
where $\theta_{a}\text{:}x^{i}$ denotes the iteration training on $x^{i}$. However, it is impractical to store model parameters and the Gaussian noise for each training iteration. Thus, for a diffusion model with parameter $\theta$, given a test generation $s$, we estimates the influence of a training data sample $x^{i}$ with respect to test generation $s$ by:
$\mathcal{I}_{\theta}(x^{s},x^{i})=e\bar{\eta}\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}\Bigl{(}f_{\theta}(z^{i}_{t},t),\epsilon\Bigr{)}\nabla_{\theta}\mathcal{L}\Bigl{(}f_{\theta}(z^{s}_{t},t),\epsilon\Bigr{)}$ (7)
where $e$ is the number of epochs, $\bar{\eta}$ is the average learning rate during training, $\epsilon$ corresponds to the Gaussian noise used in the training process. However, storing all Gaussian noise from the training process is impractical. Therefore, we randomize the Gaussian noise following the same distribution as in the training process for influence estimation.
@@
$\mathcal{I}_{\theta}(X^{s},X^{i})=e\bar{\eta}\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}\Bigl{(}f_{\theta}(z^{i}_{p},z^{i}_{t},t),\epsilon\Bigr{)}\nabla_{\theta}\mathcal{L}\Bigl{(}f_{\theta}(z^{s}_{p},z^{s}_{t},t),\epsilon\Bigr{)}$ (8)
where $p^{i}$ is the prompt of training data sample, $p^{s}$ denotes the prompt of test generation, $z^{i}_{p}$, $z^{s}_{p}$ is the embedding of prompt $p^{i}$ and $p^{s}$, respectively.
@@ 3 DMin: Scalable Influence Estimation
For a given generated image $x^{s}$ and the corresponding prompt $p^{s}$, the objective of DMin is to estimate an influence score $\mathcal{I}_{\theta}(X^{s},X^{i})$ for each training pair $X^{i}=(p^{i},x^{i})$, where $X^{s}=(p^{s},x^{s})$. Based on Equation 8, the $\mathcal{I}_{\theta}(X^{s},X^{i})$ can be expressed as the summation of the inner product between the loss gradients of the training sample and the generated image, computed with respect to the same noise $\epsilon$ across timesteps $t\in\{1,2,\cdots,T\}$. Since the training dataset is fixed and remains unchanged after training, a straightforward approach is to cache or store the gradients of each training sample across timesteps. When estimating the influence for a given query generated image, we only need to compute the gradient for the generated image and perform inner product with the cached gradients of each training sample.
@@
In this section, we explain how we reduce the storage requirements for caching such large gradients from gigabytes to kilobytes (Gradient Computation) and how we perform influence estimation for a given generated image on the fly (Influence Estimation), as shown in Figure 2. We use stable diffusion with text-to-image task as an example in this section; similar procedures can be applied to other models.
@@ 3.1 Gradient Computation
Forward and Backward Passes. In forwarding, following the same process as training, for a training pair $(p^{i},x^{i})$ and a timestep $t$, the prompt $p^{i}$ is passed through the encoder to obtain a prompt embedding, while the image $x^{i}$ is passed through a VAE to obtain a latent representation $z_{0}^{i}$. Gaussian noise is then generated from $\epsilon\sim\mathcal{N}(0,I)$ and added to the latent representation to obtain a noisy latent presentation. Then input the timestep $t$, noisy latent presentation $z^{i}_{t}=z_{0}^{i}+\epsilon$, and the embedding $z_{p}^{i}$ are then fed into the model for the forward pass.
@@
Inspired by the prior works on vector compression [21, 22], we compress the gradient vector through four steps: (1) padding, (2) permutation, (3) random projection, and (4) group addition. In the gradient compression process, we first pad the gradient vector to the smallest length $L_{\text{pad}}$ that can be evenly divided by $v$. Padding can be achieved by appending 0s to the original gradient vector until the desired length is reached. Next, we permute the gradient vector using a random permutation to disrupt any inherent structure in the vector representation. We then perform an element-wise multiplication of the permuted gradient vector with a random projection vector. The random projection vector is of the same length as the gradient vector and consists of elements randomly set to either -1 or 1 with equal probability. This step projects the gradient onto a randomized basis, reducing redundancy while preserving essential information. Finally, we divide the $L_{\text{pad}}$ elements of the gradient vector into $\frac{L_{\text{pad}}}{v}$ groups, summing up the elements within each group to produce the compressed vector of dimension $v$.
@@
Normalization. Some prior studies have highlighted the inherent instability of gradients in deep learning [22, 2, 9, 12] particularly in extremely large models. This instability arises from the potential for unusually large weights and gradients in the model. In our experiments, we encountered this issue: the magnitude of some gradient values is found to be extremely large. Such large gradient values can dominate the inner product, leading to incorrect results. To address this, we apply L2 normalization to the gradient vector before compression, which effectively mitigates the impact of unusually large gradient magnitudes. Consequently, Equation 8 can be reformulated as:
$\mathcal{I}_{\theta}(X^{s},X^{i})=$ (9)

### v4

@@ Abstract
Identifying the training data samples that most influence a generated image is a critical task in understanding diffusion models (DMs), yet existing influence estimation methods are constrained to small-scale or LoRA-tuned models due to computational limitations. To address this challenge, we propose DMin (Diffusion Model influence), a scalable framework for estimating the influence of each training data sample on a given generated image. To the best of our knowledge, it is the first method capable of influence estimation for DMs with billions of parameters. Leveraging efficient gradient compression, DMin reduces storage requirements from hundreds of TBs to mere MBs or even KBs, and retrieves the top-$k$ most influential training samples in under 1 second, all while maintaining performance. Our empirical results demonstrate that DMin is both effective in identifying influential training samples and efficient in terms of computational and storage requirements.
@@ 1 Introduction
Diffusion models have emerged as powerful generative models, capable of producing high-quality images and media across various applications [8, 41, 25, 43, 22]. Despite their impressive performance, the datasets used for training are often sourced broadly from the internet [38, 40, 39, 31, 23]. This vast dataset diversity allows diffusion models to generate an extensive range of content, enhancing their versatility and adaptability across multiple domains [19, 5]. However, it also means that these models may inadvertently generate unexpected or even harmful content, reflecting biases or inaccuracies present in the training data.
@@
Recently, many studies have explored influence estimation in diffusion models [28, 18, 28, 30, 11]. These methods assign an influence score to each training data sample relative to a generated image, quantifying the extent to which each sample impacts the generation process. For instance, DataInf [18] and K-FAC [28] are influence approximation techniques tailored for diffusion models. However, they are both second-order methods that require the inversion of the Hessian matrix. To approximate this inversion, they must load all the gradients of training data samples across several predefined timesteps. Notably, in the case of the full-precision Stable Diffusion 3 Medium model [10], the gradient of the entire model requires approximately 8 GB of storage. Collecting gradients for one training sample over 10 timesteps would consume $8\times 10=80$ GB. Scaling this requirement to a training dataset of $10,000$ samples results in a storage demand of around 800 TB – far exceeding the capacity of typical memory or even hard drives. Given that diffusion models are often trained on datasets with millions of samples, this storage demand becomes impractical. Consequently, these methods are limited to LoRA-tuned models or small diffusion models [15, 36]. Although some prior works have applied gradient compression, such as SVD [13] and quantization [28], the achieved compression rates are insufficient to maintain performance at this scale.
@@
Alternatively, Journey-TRAK [11] and D-TRAK [30] are first-order methods for influence estimation on diffusion models, which are extended from TRAK [32] on deep learning models. Both approaches utilize random projection to reduce the dimensionality of gradients. However, for large diffusion models, such as the full-precision Stable Diffusion 3 Medium model, the gradient dimensionality exceeds 2 billion parameters. Using the suggested projection dimension of $32,768$ in D-TRAK, storing such a $2\text{B}\times 32,768$ projection matrix requires more than 238 TB of storage. Even if the projection matrix is dynamically generated during computation, the scale of these operations substantially slows down the overall process. As a result, they are only feasible for small models or adapter-tuned models.
@@
Challenges. Although these approaches have demonstrated superior performance on certain diffusion models, several key challenges remain: (1) Scalability on Model Size: Existing methods either require computing second-order Hessian inversion or handling a massive projection matrix, both of which restrict their applicability to large diffusion models. (2) Scalability on Dataset Size: Diffusion models frequently rely on datasets containing millions of samples, making the computation of a Hessian inversion for the entire training dataset impractical. Additionally, storing the full gradients for all training data samples presents a significant challenge. (3) Fragility of Influence Estimation: Previous studies have demonstrated the fragility of influence estimation in extremely deep models [24, 2, 9, 12]. Similarly, we observed this fragility in large diffusion models, regardless of whether they use U-Net or transformer.
@@
  • Our experimental results demonstrate DMin’s effectiveness and efficiency in influence estimation.
  • We provide an open-source PyTorch implementation with multiprocessing support [^1] We will release the GitHub link after acceptance..
@@ 2 Influence Estimation for Diffusion Models
For a latent diffusion model, data $x_{0}$ is first encoded into a latent representation $z_{0}$ using an encoder $E$ by $z_{0}=E(x_{0})$. The model then operates on $z_{0}$ through a diffusion process to introduce Gaussian noise and iteratively denoise it. The objective is to learn to reconstruct $z_{0}$ from a noisy latent $z_{t}$ at any timestep $t\in\{1,2,\cdots,T\}$ in the diffusion process, where $T$ is the number of diffusion steps. Let $\epsilon_{t}\sim\mathcal{N}(0,I)$ denote the Gaussian noise added at timestep $t$. We define the training objective at each timestep $t$ as follows:
@@
where $\theta$ represents the model parameters, $z_{t}$ is the noisy latent representation of $z_{0}$ at timestep $t$, and $f_{\theta}(z_{t},t)$ represents the model’s predicted noise at timestep $t$ for the noisy latent $z_{t}$. $\mathcal{L}(\cdot)$ is the loss function between the predicted noise and the actual Gaussian noise.
Given a test generation $x^{s}$, where $x^{s}$ is generated by a well-trained diffusion model with parameters $\theta^{*}$, the goal of influence estimation is to estimate the influence of each training data sample $x^{i}$ ($1\leq i\leq N$) on generating $x^{s}$, where NN is the size of the training dataset. Let $z^{i}_{t}$ represent the latent representation of $x^{i}$ at timestep $t$, and let $z^{s}_{t}$ denote the latent representation of the test generation $x^{s}$ at timestep $t$.
In the $\alpha$-th training iteration, the model parameters $\theta_{\alpha+1}$ are updated from $\theta_{\alpha}$ by gradient descent on the noise prediction loss for batch $B=(B_{z},B_{t},B_{\epsilon})$:
@@
where $\eta_{\alpha}$ denotes the learning rate in the $\alpha$-th iteration, $(z_{t}^{i},t^{i},\epsilon_{t}^{i})\in B$, and the contribution of $(z_{t}^{i},t^{i},\epsilon_{t}^{i})$ to the batch gradient is $\frac{1}{|B|}\nabla_{\theta_{\alpha}}\mathcal{L}(f_{\theta_{\alpha}}(z^{i}_{t},t^{i}),\epsilon^{i}_{t})$. The influence of this training iteration $z^{i}_{t}$ with respect to $z^{s}_{t}$ on timestep $t$ can be quantified as the change in loss:
@@
where $z^{s}_{t}=E(x^{s})+\epsilon^{i}_{t}$, denoting the latent representation of $x^{s}$ after adding Gaussian noise $\epsilon^{i}_{t}$, and $\mathcal{I}_{\theta_{\alpha+1},t}(x^{s},x^{i})$ represents the influence of $x^{i}$ with respect to $x^{s}$ at the $\alpha$-th iteration with timestep $t$. Then $\mathcal{L}\Bigl(f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr)$ can be expanded via Taylor expansion:
$\hskip-3.61371pt\mathcal{L}\Bigl(f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr)=\mathcal{L}\Bigl(f_{\theta_{\alpha}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr)$ (4)
[…]
$\hskip-3.61371pt\mathcal{I}_{\theta_{\alpha+1},t}(x^{s},x^{i})$ $=\mathcal{L}\Bigl(f_{\theta_{\alpha+1}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr)-\mathcal{L}\Bigl(f_{\theta_{\alpha}}(z^{s}_{t},t),\epsilon^{i}_{t}\Bigr)$ (5)
@@
To estimate the influence of training data sample $x^{i}$, we sum over all training iterations that use $x^{i}$ and over timesteps $t\in\{1,2,\cdots,T\}$:
$\hskip-7.22743pt\mathcal{I}_{\theta^{*}}(x^{s},x^{i})=$ (6)
@@
where $\theta_{a}\text{:}x^{i}$ denotes training iterations that use $x^{i}$. However, it is impractical to store model parameters and the Gaussian noise for each training iteration. Thus, for a diffusion model with parameters $\theta$, given a test generation $x^{s}$, we estimate the influence of a training data sample $x^{i}$ with respect to $x^{s}$ by:
$\hskip-7.22743pt\mathcal{I}_{\theta}(x^{s},x^{i})=e\bar{\eta}\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}\Bigl(f_{\theta}(z^{i}_{t},t),\epsilon\Bigr)\nabla_{\theta}\mathcal{L}\Bigl(f_{\theta}(z^{s}_{t},t),\epsilon\Bigr)$ (7)
where $e$ is the number of epochs, $\bar{\eta}$ is the average learning rate during training, $\epsilon$ corresponds to the Gaussian noise used in the training process. However, storing all Gaussian noise from the training process is impractical. Therefore, for influence estimation we draw Gaussian noise from the same distribution as in training.
@@
$\hskip-7.22743pt\mathcal{I}_{\theta}(X^{s},X^{i})=e\bar{\eta}\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}\Bigl(f_{\theta}(z^{i}_{p},z^{i}_{t},t),\epsilon\Bigr)\nabla_{\theta}\mathcal{L}\Bigl(f_{\theta}(z^{s}_{p},z^{s}_{t},t),\epsilon\Bigr)$ (8)
where $p^{i}$ is the prompt of the training data sample, $p^{s}$ denotes the prompt of the test generation, and $z^{i}_{p}$ and $z^{s}_{p}$ are the embeddings of prompts $p^{i}$ and $p^{s}$, respectively.
The summation over timesteps in Equation 8 should be interpreted as a first-order approximation of the training trajectory, rather than an assumption that timestep contributions are statistically independent. In this formulation, cross-timestep dependencies are implicitly reflected in the final parameter state $\theta$ where all gradients are evaluated. For scalability, we omit second-order cross-timestep terms, consistent with first-order influence estimation methods.
@@ 3 DMin: Scalable Influence Estimation
For a given generated image $x^{s}$ and the corresponding prompt $p^{s}$, the objective of DMin is to estimate an influence score $\mathcal{I}_{\theta}(X^{s},X^{i})$ for each training pair $X^{i}=(p^{i},x^{i})$, where $X^{s}=(p^{s},x^{s})$. Based on Equation 8, $\mathcal{I}_{\theta}(X^{s},X^{i})$ can be expressed as the sum of inner products between the loss gradients of the training sample and the generated image, computed with respect to the same noise $\epsilon$ across timesteps $t\in\{1,2,\cdots,T\}$. Since the training dataset is fixed and remains unchanged after training, a straightforward approach is to cache or store the gradients of each training sample across timesteps. When estimating the influence for a given query generated image, we only need to compute the gradient for the generated image and perform inner product with the cached gradients of each training sample.
@@
In this section, we explain how we reduce the storage requirements for caching such large gradients from gigabytes to kilobytes (Gradient Computation) and how we perform influence estimation for a given generated image on the fly (Influence Estimation), as shown in Figure 2. We use Stable Diffusion on the text-to-image task as an example in this section; similar procedures can be applied to other models.
@@ 3.1 Gradient Computation
Forward and Backward Passes. In the forward pass, following the same process as training, for a training pair $(p^{i},x^{i})$ and a timestep $t$, the prompt $p^{i}$ is passed through the encoder to obtain a prompt embedding, while the image $x^{i}$ is passed through a VAE to obtain a latent representation $z_{0}^{i}$. Gaussian noise is then sampled from $\epsilon\sim\mathcal{N}(0,I)$ and added to the latent representation to obtain a noisy latent representation. The timestep $t$, the noisy latent $z^{i}_{t}=z_{0}^{i}+\epsilon$, and the embedding $z_{p}^{i}$ are then fed into the model for the forward pass.
@@
Inspired by prior work on vector compression [21, 24], we compress the gradient vector through four steps: (1) padding, (2) permutation, (3) random projection, and (4) group addition. In the gradient compression process, we first pad the gradient vector to the smallest length $L_{\text{pad}}$ that can be evenly divided by $v$. Padding can be achieved by appending 0s to the original gradient vector until the desired length is reached. Next, we permute the gradient vector using a random permutation to disrupt any inherent structure in the vector representation. We then perform an element-wise multiplication of the permuted gradient vector with a random projection vector. The random projection vector is of the same length as the gradient vector and consists of elements randomly set to either -1 or 1 with equal probability. This step projects the gradient onto a randomized basis, reducing redundancy while preserving essential information. Finally, we divide the $L_{\text{pad}}$ elements of the gradient vector into $\frac{L_{\text{pad}}}{v}$ groups, summing up the elements within each group to produce the compressed vector of dimension $v$.
@@
Normalization. Some prior studies have highlighted the inherent instability of gradients in deep learning [24, 2, 9, 12], particularly in extremely large models. This instability arises from the potential for unusually large weights and gradients in the model. In our experiments, we encountered this issue: the magnitude of some gradient values is found to be extremely large. Such large gradient values can dominate the inner product, leading to incorrect results. To address this, we apply L2 normalization to the gradient vector before compression, which effectively mitigates the impact of unusually large gradient magnitudes. Consequently, Equation 8 can be reformulated as:
$\hskip-3.61371pt\mathcal{I}_{\theta}(X^{s},X^{i})=$ (9)